Notice that the remote program’s STDOUT shows up the local computer. The output clearly shows that the command ran on the remote computer, not the local one. Now let’s use a program requiring both STDOUT and STDIN: UMENU. On the remote computer, UMENU is run with the command:

[slitt@mydesk rjust]$
slitt@mylap2's password:
         UMENU 0.7.1 menu system (ruby version), (C) 2005 by Steve Litt
                            NO WARRANTY! See GPL license.
*****  Main Menu  *****        [s]
     Z: Zuname
...  I: Internet
...  O: Outlining
...  C: Calendars
...  N: coNfiguration
...  S: Sales
...  P: Projects
...  A: Accounting
...  L: Litt's Pet Menu
...  E: Entertainment Menu
...  W: Windows Apps (Crossover Office)
     K: Kt to MPH converter
...  H: Hurricanes
...  T: Timekeeping
     X: eXit Your choice please==>
sh: /dev/tty: No such device or address
|  Copyright (C) 2005 by Steve Litt (slitt@troubleshooters.com)            |
|  GNU General Public License version 2. NO WARRANTY!                      |
|                                                                          |
|  You should have received a copy of the GNU General Public License       |
|  along with this program; if not, write to the Free Software             |
|  Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.               |
|                                                     |
Thank You for using the Universal Menu System!
[slitt@mydesk rjust]$
0
[slitt@mydesk rjust]$

In the preceding experiment, I ran UMENU in a terminal on I know it was run on , because